OpenStudy (anonymous):

2cos4x+2sin^2 2x+4cos^2 2x=6

OpenStudy (anonymous):

2cos(4x)+2(1-cos^2(2x))+4(cos(4x)-1/2)=6 2cos(4x)+2(1-((cos(4x)-1)/2))+4((cos(4x)-1)/2) = 6

OpenStudy (anonymous):

it's enough ??

OpenStudy (anonymous):

well i got somehow cos4x=3 , but i think it should be between [-1,1]

OpenStudy (anonymous):

oh i got cos 4x= 5/3 but i still think it's wrong!

OpenStudy (anonymous):

i think the right value is x=0

OpenStudy (anonymous):

well give me a minute or two :)

OpenStudy (anonymous):

i find 3cos(4x)=3

OpenStudy (anonymous):

i use cos^2(x)+sin^2(x)=1 and cos(4x) = 2cos^2(x)-1

OpenStudy (anonymous):

but can you show me step after that second you told me, because cos4x and stuff like that is totally new for me!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

2cos(4x)+2(1-cos^2(2x))+4cos(2x) = 6 2cos(4x) +2-2cos^2(2x)+4cos(2x) =6 2cos(4x)+2cos^2(2x)=4 --------------- because of cos(4x) = 2cos^2(x)-1 then 2cos^2(2x)=cos(4x)+1 ----------------------- 2cos(4x)+cos(4x)+1=4 3cos(4x)=3 x= kpi/2

OpenStudy (anonymous):

thanks a lot !

OpenStudy (anonymous):

i hope i helped you

OpenStudy (anonymous):

well that helps alot :)

OpenStudy (anonymous):

btw. in start there was ...4cos^2(2x)=6 and in your start there is ....4cos(2x)=6, why is that

OpenStudy (anonymous):

just i forget the square 2

OpenStudy (anonymous):

but , it stays right (i forget to write it just)